# Ladder Symmetry: The Necessary and Sufficient Condition for Vanishing Love Numbers

### v1

@@ Acknowledgement
We would like to express our heartfelt gratitude to Soham Acharya, Rajes Ghosh, and Sumanta Chakraborty for their valuable discussions and helpful inputs. We also thank Lam Hui, Luca Santoni, and Alessandro Podo for their useful comments on the results during the “Theoretical Tools for Gravitational Waves Physics” workshop organized in ETH Zurich, where the results were presented by SR. SS’s research is supported by the Department of Science and Technology, Government of India, under the ANRF CRG Grant (No. CRG/2023/000934).
@@
Appendix A Form of $f_{0}(r)$ for the restricted class of KRZ metrics from [Sharma:2024hlz]
@@ Appendix B Product of Sums to Sum of Products: A closed form
In this appendix, we’ll demonstrate how our method of converting the infinite product into an infinite summation helps us cast the effective potential into a parametrized form as in [Cardoso:2019mqo], and how it should help us in detecting deviations from Ladder symmetry.
@@ Appendix D Bases of non-running TLNs
If one were to assume $\gamma_{2\ell+1}=0$, then it can be easily seen that for $2\ell+1\leq n\leq j-3$, $\gamma_{n}=0$. However, this leads to having non-zero values for $\gamma_{n}$’s with $j-2\leq n<\infty$ owing to the non-zero contributions from the $\phi_{0}^{(n)}$’s. Upon performing the infinite summation over $\gamma_{n}(r_{+}/r)^{n}$, this leads to a potentially singular function at the horizon. In fact, this is related to the choice of arbitrary constants while directly solving for $\phi_{1}(r)$ from its differential equation, and the $\gamma_{2\ell+1}=0$ choice represents a bad choice of these constants that doesn’t agree with the boundary condition of having a regular function at $r=r_{+}$, giving a logarithmic divergence there (as has been seen for different combinations of $(\ell,j)$ values.) This choice of $\gamma_{2\ell+1}=0$ also doesn’t agree with its corresponding value at the limit $r_{-}\to 0$ as mentioned in [Katagiri:2023umb], i.e., for the case where the radius of the inner horizon vanishes perturbatively. For all these reasons, $\gamma_{2\ell+1}$ must be non-zero.
Following the results of [Sharma:2024hlz], since the existence of Ladder symmetry in static, spherically symmetric metrics guarantees the vanishing of Love numbers for all $\ell$ values, having a non-zero Love number for any value of $\ell$ provides a good enough counter-example for the breaking of Ladder symmetry in such spacetimes. In the following table (Table I), we have listed the values of static tidal Love numbers for several $(\ell,j)$ combinations satisfying $j\geq 2\ell+4$.
@@ Appendix E Parametrized Love numbers for KRZ class of Metric
To obtain the love number at this order, we expand the solution at large r and compare it with equation $(17)$ of [Katagiri:2023umb]. We get

### v2

@@ Acknowledgement
We would like to express our heartfelt gratitude to Soham Acharya, Rajes Ghosh, Sreejith Nair, and Sumanta Chakraborty for their valuable discussions and helpful inputs. We also thank Lam Hui, Luca Santoni, and Alessandro Podo for their useful comments on the results during the “Theoretical Tools for Gravitational Waves Physics" workshop organized in ETH Zurich, where the results were presented by SR. SS’s research is supported by the Department of Science and Technology, Government of India, under the ANRF CRG Grant (No. CRG/2023/000934).
@@
Appendix A Form of $f_{0}(r)$ for the restricted class of KRZ metrics from [41]
@@ Appendix B Product of Sums to Sum of Products: A closed form
In this appendix, we’ll demonstrate how our method of converting the infinite product into an infinite summation helps us cast the effective potential into a parametrized form as in [13], and how it should help us in detecting deviations from Ladder symmetry.
@@ Appendix D Bases of non-running TLNs
If one were to assume $\gamma_{2\ell+1}=0$, then it can be easily seen that for $2\ell+1\leq n\leq j-3$, $\gamma_{n}=0$. However, this leads to having non-zero values for $\gamma_{n}$’s with $j-2\leq n<\infty$ owing to the non-zero contributions from the $\phi_{0}^{(n)}$’s. Upon performing the infinite summation over $\gamma_{n}(r_{+}/r)^{n}$, this leads to a potentially singular function at the horizon. In fact, this is related to the choice of arbitrary constants while directly solving for $\phi_{1}(r)$ from its differential equation, and the $\gamma_{2\ell+1}=0$ choice represents a bad choice of these constants that doesn’t agree with the boundary condition of having a regular function at $r=r_{+}$, giving a logarithmic divergence there (as has been seen for different combinations of $(\ell,j)$ values.) This choice of $\gamma_{2\ell+1}=0$ also doesn’t agree with its corresponding value at the limit $r_{-}\to 0$ as mentioned in [25], i.e., for the case where the radius of the inner horizon vanishes perturbatively. For all these reasons, $\gamma_{2\ell+1}$ must be non-zero.
Following the results of [41], since the existence of Ladder symmetry in static, spherically symmetric metrics guarantees the vanishing of Love numbers for all $\ell$ values, having a non-zero Love number for any value of $\ell$ provides a good enough counter-example for the breaking of Ladder symmetry in such spacetimes. In the following table (Table I), we have listed the values of static tidal Love numbers for several $(\ell,j)$ combinations satisfying $j\geq 2\ell+4$.
@@ Appendix E Parametrized Love numbers for KRZ class of Metric
To obtain the love number at this order, we expand the solution at large r and compare it with equation $(17)$ of [25]. We get